OpenStudy (anonymous):

Solve and check. 7 ^-x = 21

OpenStudy (anonymous):

I think you have to use log(x), so: -x log(7)=log(21) -x = log(21)/log(7) X = -(log(21)/log(7))

OpenStudy (anonymous):

The answers are -2.687 -1.565 -3.729 -1.125

OpenStudy (anonymous):

I just dont understand how to get the answer

OpenStudy (anonymous):

just use calculator & find value of log(21) & log(7) & plug that values in last step You will get the answer

OpenStudy (anonymous):

& simplify it

OpenStudy (anonymous):

If you want to take down the -x from 7^-x, you need to logarithm it, so: 7^-x (use log) ---> -x*log (7) But if you make a change at the left side of the equal mark (=), then you have to do it on the right side too. So 21 (use log) ---> log(21)

OpenStudy (anonymous):

Do I divide?

OpenStudy (reemii):

the first answer is good. try to understand each steps. however for the answer, indeed, it's just \(-\log(21)/\log(7)\) to enter in the calculator. I obtain -1.565

OpenStudy (anonymous):

After you log each side of the equation, you want to make x by itself, so you throw over the log(7) to the other side.